# SpotVLM: Cloud-edge Collaborative Real-time VLM based on Context Transfer

### v1

@@
SpotVLM: Cloud-edge Collaborative Real-time VLM
based on Context Transfer
Chen Qian1, Xinran Yu1, Zewen Huang1, Danyang Li1, Qiang Ma1,
Fan Dang2, Xuan Ding1, Guangyong Shang3, Zheng Yang1
1 Abstract
Vision–Language Models (VLMs) are increasingly deployed in real-time applications such as autonomous driving and human–computer interaction, which demand fast and reliable responses based on accurate perception. To meet these requirements, existing systems commonly employ cloud–edge collaborative architectures, such as partitioned Large Vision-Language Models (LVLMs) or task offloading strategies between Large and Small Vision-Language Models (SVLMs). However, these methods fail to accommodate cloud latency fluctuations and overlook the full potential of delayed but accurate LVLM responses. In this work, we propose a novel cloud–edge collaborative paradigm for VLMs, termed Context Transfer, which treats the delayed outputs of LVLMs as historical context to provide real-time guidance for SVLMs inference. Based on this paradigm, we design SpotVLM, which incorporates both context replacement and visual focus modules to refine historical textual input and enhance visual grounding consistency. Extensive experiments on three real-time vision tasks across four datasets demonstrate the effectiveness of the proposed framework. The new paradigm lays the groundwork for more effective and latency-aware collaboration strategies in future VLM systems.
@@ 2 Introduction
VLMs have witnessed rapid advancements in recent years. These models typically consist of a visual encoder, a language decoder, and a cross-modal fusion module, enabling them to perform a wide range of tasks such as image captioning (Xue et al. 2025), visual question answering (Fang et al. 2025), and robotic perception (Zhou et al. 2025). Notably, LVLMs (Liu et al. 2025), empowered by large-scale model capacity and extensive training data, demonstrate superior generalization ability and reasoning performance across a broad range of downstream tasks. However, due to their substantial computational demands, these models are often deployed on the cloud servers, which inevitably involve network transmission, model inference, and request queuing delays, posing challenges for scenarios that require real-time responses within seconds.
To overcome these limitations, recent studies have proposed lightweight SVLMs for edge deployment (Yao et al. 2024; Rang et al. 2025), enabling low-latency, on-device inference suitable for real-time scenarios like autonomous driving (Lu et al. 2025; Sima et al. 2024; Tian et al. 2024), augmented reality (Pan and Liu 2025), and smart wearable devices (Huang et al. 2024). However, due to their limited model capacity, SVLMs still exhibit significantly lower inference accuracy than LVLMs (Zhao et al. 2025).
To enable efficient and accurate real-time visual reasoning, recent studies (Hu et al. 2024b; Li et al. 2025b; Hu et al. 2024a) have explored cloud–edge collaborative architectures. As illustrated in Fig.1, existing approaches fall into two categories: (i) Model partitioning, which splits the LVLM across the cloud and edge based on hardware constraints. For example, Distributed VLM (Li et al. 2025b) deploys the visual encoder on the edge and transmits extracted features to the cloud, where they are fused with text and processed by an LLM. (ii) Dynamic task offloading, which routes inputs based on task complexity and system conditions. For instance, ADAS (Hu et al. 2024b) offloads tasks to the cloud for LVLM inference by considering factors such as latency and quality of service (QoS).
Although such methods reduce cloud-side inference workload, they still face performance bottlenecks from service latency. In video streaming scenarios, accumulated delays often cause the outputs of LVLMs arriving several frames behind the current input (e.g., producing results for frame t–d while the system is processing frame t). Consequently, although these outputs are of high quality, they are temporally outdated and cannot be effectively utilized for real-time reasoning.
@@
$\bullet$ Utilization of High-Quality Textual Context. Historical context plays a crucial role in generating high-quality outputs. For example, Qwen-VL (Bai et al. 2023) adopts an accumulative input structure in multi-turn interactions: each round of input consists of the input and response from the previous round, along with the current query. This context is progressively extended until the window reaches the maximum memory or model-defined limit. Such accumulation allows the model to integrate information across turns, enhancing semantic consistency and reasoning depth. More importantly, the generated outputs are recursively fed into subsequent inputs, forming a performance-enhancing feedback loop. Although small models adopt similar input strategies, their limited generative capacity often results in lower-quality historical content, which negatively affects reasoning performance. Enhancing the quality of contextual history thus emerges as a key factor in improving the reasoning ability of small models.
$\bullet$ Grounding of Semantically Salient Image Regions. The ability to align visual content with textual descriptions is critical in vision-language reasoning. However, SVLMs often struggle in this aspect, particularly when dealing with fine-grained semantics or complex scenes. Prior studies (Zheng et al. 2025; Xu et al. 2025; Zhang et al. 2025a) have demonstrated that LVLMs, leveraging their grounding capabilities, can automatically crop semantically rich image regions as refined inputs—thereby enhancing the focus and semantic interpretability in visual reasoning. Guiding small models to focus on task-relevant image regions enhances their ability to interpret complex visual and linguistic cues, ultimately improving their reasoning performance.
Our Work: Given the inherent limitations of cloud-based LVLMs in terms of latency and response stability, we propose a key design principle in this work: Context Transfer. Instead of relying on the delayed outputs from LVLMs as real-time feedback, they are reused as contextual priors to enhance future reasoning in SVLMs.
Building on this principle, we introduce SpotVLM, which incorporates two collaborative modules that guide the SVLM along both the linguistic and visual dimensions to achieve more accurate multimodal reasoning.
Motivated by the strong impact of historical context quality on VLM inference, we design a Context Replacement Module (CRM). Specifically, It extracts key textual segments from high-quality LVLM responses, reorganizes them based on task context, and embeds them as contextual history to replace the original low-quality history. These informative exemplars implicitly guide the small model to focus on semantically relevant visual cues that were previously overlooked, thereby enhancing cross-modal association and reasoning performance.
@@
By providing high-quality textual history and guiding attention to semantically relevant visual regions, these modules effectively leverage the delayed outputs of LVLMs to compensate for the limited reasoning capacity of SVLMs.
@@
  • We propose a novel architectural design paradigm, termed Context Transfer, for real-time vision-language reasoning. Instead of relying on the final predictions from LVLMs, this paradigm treats their delayed outputs as contextual priors to guide SVLMs, enabling a new form of efficient large–small model collaboration.
  • Based on this principle, we design SpotVLM with two zero-shot collaborative modules: the Context Replacement Module and the Visual Focus Module, which guide the SVLMs to reuse high-quality outputs from the LVLMs along the linguistic and visual dimensions, respectively.
  • Extensive experiments on four public benchmarks validate the effectiveness of SpotVLM, showing consistent improvements over prior cloud–edge collaborative approaches.
@@ 3.1 Cloud-Edge Collaborative VLMs
Figure 2: Overview of the proposed collaboration framework, SpotVLM. The system takes timestamped video streams as input and performs two parallel operations: uploading selected keyframes to the cloud-based LVLM for processing, and conducting local inference using the SVLM. Given potential cloud latency, delayed LVLM outputs are reused as historical context. These outputs guide the SVLM through two modules, Context Replacement and Visual Focus, to improve the quality of real-time predictions.
Driven by the growing demand for real-time VLM-based processing, applications such as autonomous driving (Lu et al. 2025) are shifting from modular to end-to-end pipelines (Shao et al. 2024; Li et al. 2025a), where latency sensitivity is critical and excessive delays can lead to system failure.
To enable fast and robust responses, cloud-edge collaborative architectures for VLMs have attracted increasing attention. Existing collaboration paradigms can be categorized into two types. The first category is model partitioning (Li et al. 2025b; Zhang et al. 2025b). To improve processing efficiency for LVLMs in the cloud, Distributed VLM (Li et al. 2025b) offload the visual encoder to the edge device, which encodes input images rapidly and transmits the feature embeddings to the cloud for processing and decoding. VaVLM (Zhang et al. 2025b) further reduces transmission overhead by uploading only the region-of-interest (ROI) of images. This parallel architecture entirely rely on stable and responsive cloud services, making it difficult to guarantee robust frame-wise inference in real-time settings. The second category involves dynamic task allocation to different models according to predefined offloading strategies. Some systems (Hu et al. 2024b) formulate the offloading decision as an optimization problem, incorporating multiple system-level factors such as Quality of Service (QoS) and latency. In addition, frameworks such as LAECIPS (Hu et al. 2024a) pretrain a dedicated task classifier to estimate task complexity and offload difficult samples to cloud-based large models. While these approaches leverage the complementary strengths of large and small models to reduce cloud-side computation, they overlook the continued guidance that high-quality LVLM responses can provide for subsequent inference. Recognizing this underexplored potential, we propose a real-time collaborative framework that integrates delayed LVLM outputs into SVLM context to enable continual refinement and robust inference.
@@ 3.2 Historical Context in VLMs
Context refers to the sequence of prior interactions retained to guide subsequent inferences, essential for maintaining coherence in multi-turn tasks. In VLMs, this mechanism is adapted from Large Language Models (LLMs), which use fixed-length context windows to preserve relevant token sequences and enhance reasoning consistency. Prior research on LLMs has shown that adjusting context improves model performance, as reflected in translation accuracy (Sung et al. 2024) and dialogue consistency (Wei et al. 2024). Similarly, VLMs leverage effective context preservation to significantly improve robotic navigation in embodied AI scenarios (Habibpour and Afghah 2025), reducing the risk of decision paralysis and oscillatory behavior. Building on this advantage, the proposed SpotVLM retain the output from LVLMs as high-quality context to guide the SVLMs, helping it focus on salient content and improve inference quality.
@@ 4 Architecture
Real-time vision-language reasoning tasks pose dual challenges in terms of processing speed and inference accuracy. To achieve robust and high-precision outputs, this work leverages the strengths of LVLMs in textual understanding and visual perception to assist SVLMs in real time. This section illustrates the proposed collaborative architecture using a multi-object recognition task in autonomous driving at a frame rate of 1 FPS as a representative example.
@@ 4.1 Workflow
The front-facing camera continuously captures frames and sends them to the edge device. For each timestamped frame, the edge decides whether to upload it to the cloud based on the latency of the previous cloud interaction. Each frame is also processed locally by the SVLM for real-time inference. For uploaded frames, LVLM responses are handled according to latency: results within 1 second are adopted and stored, while delayed outputs are retained for later use, and real-time results from the SVLM are adopted for the current frame.
As illustrated in Fig.2, the output returned by the LVLM contains two key components: (1) the response content to the given query, denoted as <answer>; and (2) the attended image regions during inference, indicated as <ROI>. These two outputs provide guidance for subsequent frames from textual and visual perspectives, respectively. In addition, the returned result includes the timestamp of the original frame, which allows the system to assess current cloud latency and determine the temporal relevance of the result for semantic guidance.
@@ 4.2 Context Replacement
In multi-turn vision-language reasoning tasks, VLMs typically take an image and a textual question from the user as the current input. To support continuous multimodal understanding, VLMs often retain part of the historical context during inference. This insight informs the design of two novel mechanisms for enabling textual-level collaboration across LVLMs and SVLMs.
As detailed in Sec.4.1, high‑quality outputs from the large model are continuously updated and stored in an edge‑side cache as $\mathcal{H}_{T-d}^{LVLM}$:
$\mathcal{H}_{T-d}^{LVLM}\leftarrow LVLM(I_{T-d},Q_{T-d}),$ (1)
where $I_{T-d}$ and $Q_{T-d}$ denote the image and question for Frame T-d, respectively. When processing frame T, the system compares the current timestamp with the most recent timestamp T-d associated with a returned LVLM result. If the delay is within a predefined threshold $\delta$, the SVLM replaces its original output for frame T-d with the response from LVLM, and incorporates it into the historical context to guide inference on frame T.
$if\quad\Delta t=T-(T-d)\leq\delta,\quad\mathcal{H}_{T-d}^{SVLM}\leftarrow\mathcal{H}_{T-d}^{LVLM}$ (2)
Given the temporal continuity and strong correlation between adjacent frames, providing more accurate historical information increases the likelihood of correctly activating relevant tokens and suppressing irrelevant ones, thereby improving the overall precision of outputs generated by the SVLM.
Moreover, instead of directly concatenating raw questions and answers as historical context, we design an in-context learning prompt template to compress and summarize dialogue history for the SVLM, and the result is denoted as follows:
$\tilde{\mathcal{H}}=Compress(\mathcal{H}_{\leq T-1}).$ (3)
@@
Given the limited memory capacity of edge devices and the context window constraints of lightweight models, naively stacking multi-turn inputs may lead to token truncation. By compressing the historical content, our method not only reduces the number of irrelevant tokens but also enhances attention to the semantics of prior answers for the model.
@@
To summarize, with the textual collaboration strategy in place, the SVLM receives an input structure at frame T that corresponds to the right side of the Context Replacement module in the figure.
$Input_{T}^{SVLM}=\{I_{T},Q_{T},\tilde{\mathcal{H}}\}$ (4)
This input incorporates distilled historical answers from the LVLM, enabling the SVLM to perform accurate and efficient semantic reasoning.
Figure 3: ROI-based region tracking via cosine similarity. Given ROIs of traffic cone and pedestrian from frame T–d, cosine similarity highlights related regions in frame T to guide the attention of the small model.

### v2

# edgeVLM: Cloud-edge Collaborative Real-time VLM based on Context Transfer

@@
Abstract
Vision–Language Models (VLMs) are increasingly deployed in real-time applications such as autonomous driving and human–computer interaction, which demand fast and reliable responses based on accurate perception. To meet these requirements, existing systems commonly employ cloud–edge collaborative architectures, such as partitioned Large Vision-Language Models (LVLMs) or task offloading strategies between Large and Small Vision-Language Models (SVLMs). However, these methods fail to accommodate cloud latency fluctuations and overlook the full potential of delayed but accurate LVLM responses. In this work, we propose a novel cloud–edge collaborative paradigm for VLMs, termed Context Transfer, which treats the delayed outputs of LVLMs as historical context to provide real-time guidance for SVLMs inference. Based on this paradigm, we design edgeVLM, which incorporates both context replacement and visual focus modules to refine historical textual input and enhance visual grounding consistency. Extensive experiments on three real-time vision-language reasoning tasks across four datasets demonstrate the effectiveness of the proposed framework. The new paradigm lays the groundwork for more effective and latency-aware collaboration strategies in future VLM systems. Code will be publicly released before publication.
@@ 1 Introduction
VLMs have witnessed rapid advancements in recent years. These models typically consist of a visual encoder, a language decoder, and a cross-modal fusion module, enabling them to perform a wide range of tasks such as image captioning [xue2025progress], visual question answering [Fang_2025_CVPR], and robotic perception [zhou2025physvlm, liu2025reasongrounder]. Notably, LVLMs [liu2025making], empowered by large-scale model capacity and extensive training data, demonstrate superior generalization ability and reasoning performance across a broad range of downstream tasks. However, due to their substantial computational demands, these models are often deployed on the cloud servers, as illustrated in Fig.1(b). This deployment inevitably involve network transmission, model inference, and request queuing delays, posing challenges for scenarios that require real-time responses within seconds.
To overcome these limitations, recent studies have proposed lightweight SVLMs for edge deployment [yao2024minicpm, rang2025eve, lu2025bluelm], enabling low-latency [vasu2025fastvlm], on-device inference suitable for real-time scenarios like autonomous driving [lu2025can, sima2024drivelm, tian2024drivevlm], augmented reality [pan2025metaspatial], and smart wearable devices [huang2024vinci], as illustrated in Fig.1(a). However, due to their limited model capacity, SVLMs still exhibit significantly lower inference accuracy than LVLMs [zhao2025stitch].
To enable efficient and accurate real-time vision-language reasoning, recent studies [hu2024cloud, li2025distributed, hu2024laecips] have explored cloud–edge collaborative architectures. Existing approaches fall into two categories: (i) Model partitioning , which splits the LVLM across the cloud and edge based on hardware constraints. A representative example is Distributed VLM [li2025distributed], depicted in Fig.1(c), which deploys the visual encoder on the edge and transmits extracted features to the cloud. (ii) Dynamic task offloading, which routes inputs based on task complexity and system conditions, as illustrated in Fig.1(d). For instance, ADAS [hu2024cloud] offloads tasks to the cloud for LVLM inference by considering factors such as latency and quality of service (QoS). However, these approaches inherently depend on obtaining timely feedback from the LVLM to produce real-time outputs for the current frame, and they largely ignore the additional latency caused by network jitter and request queuing in cloud services.
@@
$\bullet$ Utilization of High-Quality Textual Context. Historical context plays a crucial role in generating high-quality outputs. Qwen-VL [bai2023qwen], as a representative VLM, adopts an accumulative input structure in multi-turn interactions: each round of input consists of the input and response from the previous round, along with the current query. This context is progressively extended until the window reaches the maximum memory or model-defined limit. Such accumulation allows the model to integrate information across turns, enhancing semantic consistency and reasoning depth. Due to its limited generative capacity, SVLM often produces low-quality or even erroneous historical content, which propagates to subsequent steps and degrades its reasoning performance. Enhancing the quality of contextual history thus emerges as a key factor in improving the reasoning ability of small models.
$\bullet$ Grounding of Semantically Salient Image Regions. The ability to align visual content with textual descriptions is critical in vision-language reasoning [zong2025ground]. However, SVLMs often struggle in this aspect, particularly when dealing with fine-grained semantics or complex scenes. Prior studies [zheng2025deepeyes, xu2025visual, zhang2025chain] have demonstrated that LVLMs, leveraging their grounding capabilities, can automatically crop semantically rich image regions as refined inputs, thereby enhancing the focus and semantic interpretability in visual reasoning. Guiding small models to focus on task-relevant image regions enhances their ability to interpret complex visual and linguistic cues, ultimately improving their reasoning performance.
Our Work: Motivated by the critical role of accurate and task-relevant context in model reasoning, we propose a key design principle in this work: Context Transfer. Instead of relying on the delayed outputs from LVLMs as real-time feedback, they are reused as contextual priors to enhance future reasoning in SVLMs, as illustrated in Fig.1(e).
Building on this principle, we introduce edgeVLM, which incorporates two collaborative modules that guide the SVLM along both the linguistic and visual dimensions to achieve more accurate multimodal reasoning.
Motivated by the strong impact of historical context quality on VLM inference, we design a Context Replacement Module (CRM). Specifically, It extracts key textual segments from high-quality LVLM responses, reorganizes them based on task context, and embeds them as contextual history to replace the original low-quality history. By exploiting inter-frame correlations among temporally adjacent frames, these refined historical cues offer precise contextual guidance for current-frame inference, implicitly steering the small model toward semantically relevant visual evidence that was previously overlooked.
@@
  • We propose Context Transfer, a cloud–edge collaborative paradigm for real-time VLM inference. Instead of relying on the final predictions from LVLMs, this paradigm treats their delayed outputs as contextual priors to guide SVLMs, enabling a new form of efficient large–small model collaboration.
  • Based on this principle, we design edgeVLM with two training-free collaborative modules: the Context Replacement Module and the Visual Focus Module, which guide the SVLMs to reuse high-quality outputs from the LVLMs along the linguistic and visual dimensions, respectively.
  • Extensive experiments on four public benchmarks validate the effectiveness of edgeVLM, showing consistent improvements over prior cloud–edge collaborative approaches.
Figure 2: Overview of the proposed collaboration framework, edgeVLM. The system takes timestamped video streams as input and performs two parallel operations: uploading selected keyframes to the cloud-based LVLM for processing, and conducting local inference using the SVLM. Given potential cloud latency, delayed LVLM outputs are reused as historical context. These outputs guide the SVLM through two modules, Context Replacement and Visual Focus, to improve the quality of real-time predictions.
@@ 2.1 Cloud-Edge Collaborative VLMs
Driven by the growing demand for real-time VLM-based processing, applications such as autonomous driving [lu2025can] are shifting from modular to end-to-end pipelines [shao2024lmdrive, li2025generative], where latency sensitivity is critical and excessive delays can lead to system failure.
To enable fast and robust responses, cloud-edge collaborative architectures for VLMs have attracted increasing attention. Existing collaboration paradigms can be categorized into two types. The first category is model partitioning [li2025distributed, zhang2025vavlm]. To improve processing efficiency, Distributed VLM [li2025distributed] offload the visual encoder to the edge device, which encodes input images and transmits features to the cloud for processing and decoding. VaVLM [zhang2025vavlm] reduces transmission overhead by uploading only the region-of-interest (ROI) of images. These parallel architecture entirely rely on stable and responsive cloud services, making it difficult to guarantee robust frame-wise inference in real-time settings. The second category involves dynamic task allocation to different models according to predefined offloading strategies. Some systems [hu2024cloud] formulate the offloading decision as an optimization problem, incorporating multiple system-level factors such as Quality of Service (QoS) and latency. Frameworks such as LAECIPS [hu2024laecips] pretrain a dedicated task classifier to estimate task complexity and offload difficult samples to cloud-based large models. While these approaches leverage the complementary strengths of large and small models to reduce cloud-side computation, they overlook the continued guidance that high-quality LVLM responses can provide for subsequent inference. Recognizing this underexplored potential, we propose a real-time collaborative framework that integrates delayed LVLM outputs into SVLM context to enable continual refinement and robust inference.
@@ 2.2 Historical Context in VLMs
Context refers to the sequence of prior interactions retained to guide subsequent inferences, essential for maintaining coherence in multi-turn tasks. In VLMs, this mechanism is adapted from Large Language Models (LLMs), which use fixed-length context windows to preserve relevant token sequences and enhance reasoning consistency. Prior research on LLMs has shown that adjusting context improves model performance, as reflected in translation accuracy [sung2024context] and dialogue consistency [wei2024hidden]. Similarly, VLMs leverage effective context preservation to significantly improve robotic navigation in embodied AI scenarios [habibpour2025history], reducing the risk of decision paralysis and oscillatory behavior. Building on this advantage, the proposed edgeVLM retain the output from LVLMs as high-quality context to guide the SVLMs, helping it focus on salient content and improve inference quality.
@@ 3 Architecture
Real-time vision-language reasoning tasks pose dual challenges in terms of processing speed and inference accuracy. To achieve robust and high-precision outputs, this work leverages the strengths of LVLMs in textual understanding and visual perception to assist SVLMs in real-time. This section illustrates the proposed collaborative architecture using a multi-object recognition task in autonomous driving as a representative example.
@@ 3.1 Workflow
The front-facing camera continuously captures frames and sends them to the edge device. For each timestamped frame, the edge decides whether to upload it to the cloud. Each frame is also processed locally by the SVLM for real-time inference. For uploaded frames, the returned LVLM responses are handled according to latency: results within $\tau$ second are adopted and stored, while delayed outputs are retained for later use, and real-time results from the SVLM are adopted for the current frame.
As illustrated in Fig.2, the output for frame TdT-d returned by the LVLM contains two key components: (1) the response content to the given query, denoted as <answer>; and (2) the attended image regions during inference, indicated as <ROI>. The two outputs respectively provide textual and visual guidance for SVLM inference at frame $T$.
@@ 3.2 Context Replacement
In multi-turn vision-language reasoning tasks, VLMs typically take an image and a textual question from the user as the current input. To support continuous multimodal understanding, VLMs often retain part of the historical context during inference. We find that replacing the history generated by SVLM with accurate historical context supplied by a delayed LVLM guides the SVLM to produce more accurate predictions for the current frame. To theoretically justify this empirical observation, consider the following sequential prediction formulation.
Let the prediction at time $t$ be
$\hat{\mathbf{y}}_{t}=f(\mathbf{x}_{t},\mathbf{h}_{t-1}),$ (1)
where $\mathbf{x}_{t}$ is the current input and $\mathbf{h}_{t-1}$ is the historical state from frame (t1)(t\!-\!1). Denote by $\mathbf{h}^{\star}_{t-1}$ the accurate history and by $\mathbf{h}^{\star}_{t-1}+\mathbf{e}$ an inexact history. A first-order expansion in the historical state gives
$f(\mathbf{x}_{t},\mathbf{h}^{\star}_{t-1}+\mathbf{e})\approx f(\mathbf{x}_{t},\mathbf{h}^{\star}_{t-1})+J_{h}\,\mathbf{e},$ (2)
$J_{h}:=\frac{\partial f}{\partial\mathbf{h}}\Big|_{(\mathbf{x}_{t},\mathbf{h}^{\star}_{t-1})}.$ (3)
Hence the induced output deviation is bounded by
$\big\|f(\mathbf{x}_{t},\mathbf{h}^{\star}_{t-1}+\mathbf{e})-f(\mathbf{x}_{t},\mathbf{h}^{\star}_{t-1})\big\|\leq\|J_{h}\|\,\|\mathbf{e}\|.$ (4)
Let $\ell=\big\|\cdot\big\|^{2}$, the expected excess loss at time $t$ satisfies
$\mathbb{E}[\Delta\ell]\;\lesssim\;\mathrm{tr}\!\big(J_{h}\,\Sigma_{e}\,J_{h}^{\top}\big),$ (5)
where $\Sigma_{e}$ is the covariance matrix of the historical error. Therefore, a more accurate history (smaller $\Sigma_{e}$) yields a smaller expected loss at frame $t$. This insight informs the design of two novel methods for enabling textual-level collaboration across LVLMs and SVLMs.
As detailed in Sec. 3.1, the <answer> produced by the LVLM is continuously refreshed and preserved in the edge-side cache. At frame T, the system locates in the edge cache the latest LVLM result corresponding to a prior frame T - d. If the elapsed time $d\leq\delta$, that LVLM result replaces the historical output produced by the SVLM for frame T - d and is incorporated into the context used to condition inference on frame T.
Since adjacent frames are highly correlated, higher-quality historical information more reliably steers the model toward activating relevant tokens and suppressing irrelevant ones.
@@
Given the limited memory capacity of edge devices and the context window constraints of lightweight models, naively stacking multi-turn inputs may lead to token truncation. Moreover, we design an in-context learning [kang2025context] prompt template to compress and summarize dialogue history for the SVLM.
Example. In the autonomous driving setup, the edge-side SVLM predicts “car” on the historical frame T - d. If the corresponding cloud response arrives within the staleness bound $d\leq\delta$, we replace the history with the corrected set from LVLM “traffic cone, car, pedestrian.” We then convert the Q&A-style history into a compressed snippet: “This is the image from d seconds ago; detected traffic cone, car, pedestrian.” This snippet is injected into the current input as prior.
@@
To summarize, with the textual collaboration strategy in place, the SVLM receives an input structure at frame T that corresponds to the right side of the Context Replacement module in Fig. 2. This input incorporates distilled historical answers from the LVLM, enabling the SVLM to perform accurate and efficient reasoning.